OpenStudy (anonymous):

Alright, who remembers factoring??? x^2-x-6

OpenStudy (anonymous):

(x-3)(x+2)

OpenStudy (anonymous):

(x-3)(X+2) so x= 3, -2

OpenStudy (anonymous):

could someone please explain??? i have three more questions like it...

OpenStudy (anonymous):

A "quadratic" is a polynomial that looks like this "ax2 + bx + c", where "a", "b", and "c" are just numbers. For the easy case of factoring, you will find two numbers that will not only multiply to equal the constant term "c", but also add up to equal "b"

OpenStudy (anonymous):

in this case a=1 b=-1 c=-6 so in this case -3x2= -6 (c) and adding them -3+2=-1(b)